OpenStudy (anonymous):

Simplify this expression. 3x^2(7x^2 - 9)

OpenStudy (anonymous):

\[3x ^{2}(7x^{2} - 9)\]

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

You have to work with \[7x^2 - 9\] cause \[3x^2 \] is already simple. You can think of \[7x^2-9\] as a quadratic equation. Thene, you have to know that it can be written in a factoriced form. You has to solve for x the equation \[7x^2 - 9 = 0 \] See that this is equal to \[x^2 = 9/7 \] and x = \[x = \pm \sqrt{(9/7)}\] And that is equal to \[x = \pm 3/\sqrt{7}\]

OpenStudy (anonymous):

Excuse me, do you know how to solve quadratics?

OpenStudy (anonymous):

D. But then you are not SIMPLIFING, you are just multipling....applying the distributibe propertie of the product.